The enthalpy change involved in the oxidation of glucose is -2880 KJ mol-1. 25% of this energy is available for muscular work. If 100 KJ of muscular work is needed to walk one kilometre, what is the maximum distance that a person will be able to walk after consuming 120 gm of glucose??

Grade:11

1 Answers

Arun
3 years ago
−mgh= 
100
25
​ 
 ×n(−2808KJ)
−6.25×9.8×3= 
100
25
​ 
 ×n(−2808KJ)
−183.75J= 
100
25
​ 
 ×n(−2808×10 
3
 J)
n= 
25×(−2808)(1000)
−183.75×100
​ 
 ⇒ 
25×28080
183.75
​ 
 ⇒0.00026moles.
massofglucose=n×molarmass
⇒0.00026×180.156⇒0.0468g.
​